# The Low-Temperature Phenomenology of Gap Inhomogeneity in the Cuprate Superconductors: High-Energy Granularity, Low-Energy Homogeneity, and Spectral Kinks

### v1

@@ Abstract
Scanning tunneling spectroscopy experiments on a number of cuprate superconductors have revealed that these materials are highly inhomogeneous. However, even though this inhomogeneity is well-characterized experimentally, a theoretical understanding of the effect of an inhomogeneous superconducting $d$-wave order parameter on various observables is still not complete. Here, we focus on the particular role played by the length scale of superconducting order-parameter inhomogeneity. We make use of a model involving square patches tiling the system, with each patch hosting a broadly distributed random value of the $d$-wave parameter. By using large-scale simulations, we are able to study how the size of the patches affects the correspondence between various measures of the superconducting gap and the underlying order parameter. If the length scale of the inhomogeneity is smaller than the average superconducting coherence length, the resulting $d$-wave superconductor is homogeneous. However, when the order parameter varies on the scale of the coherence length, we find the emergence of a striking low-/high-energy dichotomy, in which the low-energy regime is homogeneous while the high-energy states are strongly inhomogeneous. Kinks in the local spectra are found at the energy demarcating the homogeneous-inhomogeneous transition. We also observe that the gap extracted from the low-energy slope of the LDOS is extremely uniform. We find in both of these regimes that the distribution of the spectral gap is narrower than that of the order parameter; these start to match only when the size of the patches becomes parametrically larger than the coherence length. We comment on the applicability of these results to the cuprates, discuss the limitations of the inhomogeneous $d$-wave model, and point out where beyond-mean-field correlation effects are likely to be present in addition to inhomogeneity.
@@ I Introduction
Scanning tunneling spectroscopy (STS) has been a hugely invaluable tool for uncovering many of the mysteries held by the high-$T_{c}$ cuprates Schmidt et al. (2011); Fujita et al. (2011). The real-space resolution afforded by state-of-the-art STS measurements has facilitated numerous important discoveries about these materials, such as the presence of strong inhomogeneity in the spectral gap Pan et al. (2001); Howald et al. (2001); Lang et al. (2002); McElroy et al. (2005a, b); Fang et al. (2006); Gomes et al. (2007); Alldredge et al. (2008); Pasupathy et al. (2008); Pushp et al. (2009); Parker et al. (2010); Alldredge et al. (2013); Li et al. (2022); Tromp et al. (2023), broken-symmetry states coexisting with the superconducting state Hoffman et al. (2002a); Howald et al. (2003a, b); Hamidian et al. (2016); Edkins et al. (2019); Du et al. (2020), and quasiparticle scattering interference (QPI) Hoffman et al. (2002b); Wang and Lee (2003); Capriotti et al. (2003); McElroy et al. (2003); Zhu et al. (2004); Kohsaka et al. (2008); Fujita et al. (2014); He et al. (2014); Webb et al. (2019); Ye et al. (2024). The maturation of this technique has allowed its use beyond the low-temperature regime of the cuprates where it originally found success: successful applications of the technique to both high temperatures Gomes et al. (2007); Pasupathy et al. (2008); Pushp et al. (2009); Parker et al. (2010) and heavy overdoping Li et al. (2022); Tromp et al. (2023); Ye et al. (2024) have uncovered strong inhomogeneity even in these regimes. The story emerging from three decades of STS work on the cuprates is one in which a number of seemingly conflicting ingredients are present in some form deep inside the superconducting state: patchy inhomogeneity (once likened to “quantum salad dressing” Zaanen (2002) due to the appearance of what appears to be electronic phase separation), density-wave order, and, somehow, a sufficient amount of homogeneity needed to allow for the existence of QPI. How these three apparently mutually incompatible ingredients coexist in the cuprates has been a persistent mystery since their discovery.
@@
Low-temperature differential conductance ($dI/dV$) measurements in the superconducting cuprates generally find a number of features that are present at a wide range of dopings and even in different families of cuprates: 1) the spectral gap is patchy, with a characteristic length scale of several lattice constants Pan et al. (2001); Howald et al. (2001); Lang et al. (2002); McElroy et al. (2005a); Fang et al. (2006); Gomes et al. (2007); Alldredge et al. (2008); Pasupathy et al. (2008); Pushp et al. (2009); Parker et al. (2010); Alldredge et al. (2013); Li et al. (2022); Tromp et al. (2023); Ye et al. (2024); 2) while the spectra have the V-shaped form characteristic of $d$-wave pairing at the lowest energies, these eventually display “kinks” for large-gap regions Howald et al. (2001); McElroy et al. (2005a); Fang et al. (2006); Alldredge et al. (2008); Pushp et al. (2009); Alldredge et al. (2013); Ye et al. (2024); 3) there is a negative correlation between the spectral gap and the height of the coherence peak, so that small-gap regions display comparatively tall and sharp peaks and large-gap regions have stubbier ones Lang et al. (2002); McElroy et al. (2005a); Fang et al. (2006); Alldredge et al. (2008); Ye et al. (2024); 4) the width of the spectral-gap distribution generally scales with the mean spectral gap, so the distribution is narrow at overdoping (where the average gap is small) and wide at underdoping McElroy et al. (2005a); Alldredge et al. (2008); Gomes et al. (2007); Pasupathy et al. (2008); Pushp et al. (2009); and, perhaps most important, 5) the spectra are highly homogeneous at low bias voltage, with inhomogeneity appearing only at a higher energy scale that is set roughly by the location of the kink but is also parametrically smaller than the average spectral gap McElroy et al. (2005a); Kohsaka et al. (2008); Alldredge et al. (2008); Pushp et al. (2009); Alldredge et al. (2013). The last observation dovetails nicely with the presence of quasiparticle scattering interference at low energies which is eventually lost at an energy that matches the scale where homogeneity is lost Kohsaka et al. (2008); Fujita et al. (2014); He et al. (2014); Webb et al. (2019). These observations are qualitatively consistent with the presence of homogeneous $d$-wave superconductivity at low energies, which eventually gives way to an inhomogeneous regime at high energies characterized by broken symmetry Mukhopadhyay et al. (2019).
@@
This dichotomy between the low- and high-energy states in the underdoped cuprates has been interpreted as that arising from two distinct types of excitations. In this scheme, the low-energy states are simply the Bogoliubov quasiparticles of a homogeneous $d$-wave superconductor, while the high-energy states correspond to mysterious “pseudogap” states which break translational invariance in real space and which ostensibly originate from the parent Mott insulating state Kohsaka et al. (2008); Fujita et al. (2014); He et al. (2014); Webb et al. (2019). This explanation is a particularly compelling one since this dichotomy (in energy) is structurally similar to the dichotomy one sees in temperature in the underdoped regime: the low-temperature state is a $d$-wave superconductor, while the high-temperature state (the pseudogap) hosts translational-symmetry-breaking charge order and features the absence of any phase coherence. Consequently, the high-energy states imaged in these $dI/dV$ measurements are manifestations at low temperature of the same states that are present in the high-temperature regime above $T_{c}$. This is also consistent with the picture emerging from angle-resolved photoemission spectroscopy (ARPES), which finds evidence for two distinct gaps—one corresponding to the superconductor, and another corresponding to the pseudogap Tanaka et al. (2006); Lee et al. (2007).
@@
As compelling as this picture is, one is inevitably confronted with the question of how homogeneous $d$-wave superconductivity can be made to coexist with an exotic, spatially fluctuating pseudogap state at the lowest temperatures in the first place. This picture would appear to require that these two states be separated in energy, with something akin to a “mobility edge” separating the delocalized (“$\mathbf{k}$-space”) low-energy states from the localized (“$\mathbf{r}$-space”) high-energy states. Given that the cuprates are strongly correlated materials, it is not hard to imagine that such a possibility can arise from a very complicated microscopic model that can generate both the low-energy translationally-invariant superconductor and the high-energy translational-symmetry-breaking pseudogap simultaneously. The fact that simple paradigmatic models such as the Hubbard model do not yield such behavior easily can lead one to think that any putative model explanation for this necessarily requires a complicated mechanism underlying this separation of scales in energy. (We must mention that an important advance is the work by Lee et al., which makes use of a model of “glassy nematicity” and, within a mean-field framework where the nematic order is disordered but the pairing interaction is not, is able to reproduce to some extent some of the features of this low-/high-energy dichotomy seen in STS spectra taken from the cuprates Lee et al. (2016). However, in their work, they find the homogeneous regime to lie within all energies less than the average spectral gap and the inhomogeneous regime to reside outside the spectral gap, which is a feature that is not seen in the cuprates—in experiment, the homogeneous-inhomogeneous transition occurs at a characteristic energy scale that is smaller than the average spectral gap Alldredge et al. (2008, 2013).)
@@
In this paper, we take a different tack: we leave the question of modeling the exotic, symmetry-broken pseudogap state aside and instead study a much simpler mean-field model of $d$-wave superconductivity with strong order-parameter inhomogeneity, but without any chemical-potential or hopping disorder or coexisting order. We wish to study the extent to which the aforementioned STS phenomenology can be captured within mean-field theory by a minimal model consisting only of strong order-parameter inhomogeneity, and we pay close attention to the role played by the length scale of the inhomogeneity. We do not make any assumptions about the origin of the inhomogeneity. Instead, we make use of an exceedingly simple model of gap inhomogeneity: we divide the system into equally shaped square patches, and to each patch we assign a random value of the $d$-wave order parameter, taken from a uniform distribution. (This is similar in spirit to the patch model used by Fang et al. Fang et al. (2006) and by Alvarez and Dagotto Alvarez and Dagotto (2008). It should be said that another self-consistent, ex post facto justification for the use of this model is to see how a most artificial-looking inhomogeneity model can in fact reproduce experimental spectra without much fine-tuning, as can be seen later.) The distribution of order-parameter values is fixed such that these range from zero to twice the mean; the only parameter being varied is the linear extent $l$ of the patch. The main utility of the square-patch model is the ease with which one can tune the length scale of the inhomogeneity, as opposed to other models such as the Yukawa-like model used by Nunner et al. Nunner et al. (2005, 2006); Andersen et al. (2006). We use Green’s function-based techniques to efficiently obtain the local density of states (LDOS) for very large systems (with around $3\times 10^{5}$ sites), allowing us to calculate spectral observables and statistical correlations among these observables with much accuracy. The observables we focus on are the LDOS spectra themselves and convenient experimental proxies for the true superconducting order parameter: the spectral gap, which is the location of the coherence peak, and the low-energy gap, which is the value of the gap one can extract from the low-energy slope of the LDOS, which in a bulk $d$-wave superconductor is inversely proportional to the order parameter.
@@
We find three generic regimes, depending on how large the patch length $l$ is compared to the the average superconducting coherence length $\xi$. If $l$ is less than $\xi$, the quasiparticle response of the superconductor to this order parameter profile is remarkably spatially uniform. In this regime, the LDOS from site to site does not vary appreciably. The order-parameter proxy observables show only very weak variations, and the distributions of both of these quantities are very narrow compared to that of the underlying $d$-wave order parameter. The spectra are effectively averaged over a coherence volume $\xi^{2}$, so the underlying atomic-scale variations in the order parameter are smoothed over in the quasiparticle response. If, on the other hand, $l$ is around the same magnitude as $\xi$, strikingly different behavior is observed which matches that seen in the superconducting cuprates. The LDOS at low energies is very uniform throughout the system, and deviations from uniformity only arise after a threshold energy is exceeded, beyond which the LDOS becomes strongly spatially fluctuating. We find that the kink energy demarcating the homogeneous-inhomogeneous crossover is generally set by the smallest set of gaps in the system, something that is also seen in experiment. In this regime, the statistical correlation between the local spectral gap and the height of the coherence peak is negative, similar to what is seen in STS experiments. Finally, in the regime where $l$ is larger than $\xi$, the behavior of the LDOS at a site where the local $d$-wave order parameter has a value $\Delta$ gradually reverts to that of a bulk $d$-wave superconductor with the same value of the order parameter. In this regime, the statistical correlations among the various observables also approach that expected from a bulk superconductor, with the proxy observables becoming very closely matched to the true order parameter. The locations of the kinks get pushed down all the way to zero energy with increasing $l$.
@@
We discuss in detail some ramifications of our results to the interpretation of STS experiments. In particular, we argue that in the cuprate superconductors, the behavior of the LDOS spectra and the spectral-gap distributions throughout much of the doping range are consistent with order-parameter patch sizes being in the $l\approx\xi$ regime. As a consequence, both the low-energy homogeneous $d$-wave state and the high-energy inhomogeneous state (but not the broken-symmetry density-wave and nematic states) can be explained within the same mean-field framework. We then reexamine the standard interpretation of the spectral gap as being a direct measure of the true superconducting order parameter and point out that in the ł$\l\approx\xi$ regime relevant to the cuprates, the width of the spectral-gap distribution vastly underestimates the width of the distribution of the underlying $d$-wave order parameter. It is then entirely plausible even that deep within the superconducting state of the cuprates, there can exist patches where the order parameter is very small or even zero, but proximity coupling of these with the surrounding large-order-parameter regions results in a low-energy homogeneous state within these regions that exhibits a finite spectral gap. We also consider estimates of the so-called “nodal gap” Pushp et al. (2009) and find that much of the observed “two-gap” phenomenology at low temperatures can be explained even within the inhomogeneous $d$-wave model to a surprisingly decent extent, with the dichotomy between the low-energy and high-energy states playing a crucial role. In addition, we discuss what our results mean for estimates of the local inelastic scattering rate in underdoped cuprates from STS data Alldredge et al. (2008, 2013); we find that these estimates are likely an overestimate of the true inelastic scattering rate as these do not invoke the inhomogeneity of the pairing in the estimation process. Finally, we comment on the relation of order-parameter inhomogeneity to the hitherto unsolved phenomenon of QPI extinction Kohsaka et al. (2008); Fujita et al. (2014); He et al. (2014) and suggest that inhomogeneity is of fundamental importance to this phenomenon in the first place, although we postpone a thorough discussion of this to a future paper.

### v2

@@ Abstract
Scanning tunneling spectroscopy experiments on a number of cuprate superconductors have revealed that these materials are highly inhomogeneous. However, even though this inhomogeneity is well-characterized experimentally, a theoretical understanding of the effect of an inhomogeneous superconducting $d$-wave order parameter on various observables is still not complete. Here, we focus on the particular role played by the length scale of superconducting order-parameter inhomogeneity. We make use of a model involving square patches tiling the system, with each patch hosting a broadly distributed random value of the $d$-wave parameter. By using large-scale simulations, we are able to study how the size of the patches affects the correspondence between various measures of the superconducting gap and the underlying order parameter. If the length scale of the inhomogeneity is smaller than the average superconducting coherence length, the resulting $d$-wave superconductor is homogeneous. However, when the order parameter varies on the scale of the coherence length, we find the emergence of a striking low-energy–high-energy dichotomy, in which the low-energy regime is homogeneous while the high-energy states are strongly inhomogeneous. Kinks in the local spectra are found at the energy demarcating the homogeneous-inhomogeneous transition. We also observe that the gap extracted from the low-energy slope of the local density of states is extremely uniform. We find in both of these regimes that the distribution of the spectral gap is narrower than that of the order parameter; these start to match only when the size of the patches becomes parametrically larger than the coherence length. We comment on the applicability of these results to the cuprates, discuss the limitations of the inhomogeneous $d$-wave model, and point out where beyond-mean-field correlation effects are likely to be present in addition to inhomogeneity.
@@ I Introduction
Scanning tunneling spectroscopy (STS) has been a hugely invaluable tool for uncovering many of the mysteries held by the high-$T_{c}$ cuprates [1, 2]. The real-space resolution afforded by state-of-the-art STS measurements has facilitated numerous important discoveries about these materials, such as the presence of strong inhomogeneity in the spectral gap [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], broken-symmetry states coexisting with the superconducting state [17, 18, 19, 20, 21, 22], and quasiparticle scattering interference (QPI) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. The maturation of this technique has allowed its use beyond the low-temperature regime of the cuprates where it originally found success: successful applications of the technique to both high temperatures [9, 11, 12, 13] and heavy overdoping [15, 16, 32] have uncovered strong inhomogeneity even in these regimes. The story emerging from three decades of STS work on the cuprates is one in which a number of seemingly conflicting ingredients are present in some form deep inside the superconducting state: patchy inhomogeneity (once likened to “quantum salad dressing” [33] due to the appearance of what appears to be electronic phase separation), density-wave order, and, somehow, a sufficient amount of homogeneity needed to allow for the existence of QPI. How these three apparently mutually incompatible ingredients coexist in the cuprates has been a persistent mystery since their discovery.
@@
Low-temperature differential conductance ($dI/dV$) measurements in the superconducting cuprates generally find a number of features that are present at a wide range of dopings and even in different families of cuprates: 1) the spectral gap is patchy, with a characteristic length scale of several lattice constants [3, 4, 5, 6, 8, 9, 10, 11, 12, 13, 14, 15, 16, 32]; 2) while the spectra have the V-shaped form characteristic of $d$-wave pairing at the lowest energies, these eventually display “kinks” for large-gap regions [4, 6, 8, 10, 12, 14, 32]; 3) there is a negative correlation between the spectral gap and the height of the coherence peak, so that small-gap regions display comparatively tall and sharp peaks and large-gap regions have stubbier ones [5, 6, 8, 10, 32]; 4) the width of the spectral-gap distribution generally scales with the mean spectral gap, so the distribution is narrow at overdoping (where the average gap is small) and wide at underdoping [6, 10, 9, 11, 12]; and, perhaps most important, 5) the spectra are highly homogeneous at low bias voltage, with inhomogeneity appearing only at a higher energy scale that is set roughly by the location of the kink but is also parametrically smaller than the average spectral gap [6, 28, 10, 12, 14]. The last observation dovetails nicely with the presence of quasiparticle scattering interference at low energies which is eventually lost at an energy that matches the scale where homogeneity is lost [28, 29, 30, 31]. These observations are qualitatively consistent with the presence of homogeneous $d$-wave superconductivity at low energies, which eventually gives way to an inhomogeneous regime at high energies characterized by broken symmetry [34].
@@
This dichotomy between the low- and high-energy states in the underdoped cuprates has been interpreted as that arising from two distinct types of excitations. In this scheme, the low-energy states are simply the Bogoliubov quasiparticles of a homogeneous $d$-wave superconductor, while the high-energy states correspond to mysterious “pseudogap” states which break translational invariance in real space and which ostensibly originate from the parent Mott insulating state [28, 29, 30, 31]. This explanation is a particularly compelling one since this dichotomy (in energy) is structurally similar to the dichotomy one sees in temperature in the underdoped regime: the low-temperature state is a $d$-wave superconductor, while the high-temperature state (the pseudogap) hosts translational-symmetry-breaking charge order and features the absence of any phase coherence. Consequently, the high-energy states imaged in these $dI/dV$ measurements are manifestations at low temperature of the same states that are present in the high-temperature regime above $T_{c}$. This is also consistent with the picture emerging from angle-resolved photoemission spectroscopy (ARPES), which finds evidence for two distinct gaps—one corresponding to the superconductor, and another corresponding to the pseudogap [35, 36].
@@
As compelling as this picture is, one is inevitably confronted with the question of how homogeneous $d$-wave superconductivity can be made to coexist with an exotic, spatially fluctuating pseudogap state at the lowest temperatures in the first place. This picture would appear to require that these two states be separated in energy, with something akin to a “mobility edge” separating the delocalized ($\mathbf{k}$-space”) low-energy states from the localized ($\mathbf{r}$-space”) high-energy states. Given that the cuprates are strongly correlated materials, it is not hard to imagine that such a possibility can arise from a very complicated microscopic model that can generate both the low-energy translationally-invariant superconductor and the high-energy translational-symmetry-breaking pseudogap simultaneously. The fact that simple paradigmatic models such as the Hubbard model do not yield such behavior easily can lead one to think that any putative model explanation for this necessarily requires a complicated mechanism underlying this separation of scales in energy. (We must mention that an important advance is the work by Lee et al., which makes use of a model of “glassy nematicity” and, within a mean-field framework where the nematic order is disordered but the pairing interaction is not, is able to reproduce to some extent some of the features of this low-energy–high-energy dichotomy seen in STS spectra taken from the cuprates [37]. However, in their work, they find the homogeneous regime to lie within all energies less than the average spectral gap and the inhomogeneous regime to reside outside the spectral gap, which is a feature that is not seen in the cuprates—in experiment, the homogeneous-inhomogeneous transition occurs at a characteristic energy scale that is smaller than the average spectral gap [10, 14].)
@@
In this paper, we take a different tack: we leave the question of modeling the exotic, symmetry-broken pseudogap state aside and instead study a much simpler mean-field model of $d$-wave superconductivity with strong order-parameter inhomogeneity, but without any chemical-potential or hopping disorder or coexisting order. We wish to study the extent to which the aforementioned STS phenomenology can be captured within mean-field theory by a minimal model consisting only of strong order-parameter inhomogeneity, and we pay close attention to the role played by the length scale of the inhomogeneity. We do not make any assumptions about the origin of the inhomogeneity. Instead, we make use of an exceedingly simple model of gap inhomogeneity: we divide the system into equally shaped square patches, and to each patch we assign a random value of the $d$-wave order parameter, taken from a uniform distribution. (This is similar in spirit to the patch model used by Fang et al. [8] and by Alvarez and Dagotto [38]. It should be said that another self-consistent, ex post facto justification for the use of this model is to see how a most artificial-looking inhomogeneity model can in fact reproduce experimental spectra without much fine-tuning, as can be seen later.) The distribution of order-parameter values is fixed such that these range from zero to twice the mean; the only parameter being varied is the linear extent $l$ of the patch. The main utility of the square-patch model is the ease with which one can tune the length scale of the inhomogeneity, as opposed to other models such as the Yukawa-like model used by Nunner et al. [39, 40, 41]. We use Green’s function-based techniques to efficiently obtain the local density of states (LDOS) for very large systems (with around $3\times 10^{5}$ sites), allowing us to calculate spectral observables and statistical correlations among these observables with much accuracy. The observables we focus on are the LDOS spectra themselves and convenient experimental proxies for the true superconducting order parameter: the spectral gap, which is the location of the coherence peak, and the low-energy gap, which is the value of the gap one can extract from the low-energy slope of the LDOS, which in a bulk $d$-wave superconductor is inversely proportional to the order parameter.
@@
We find three generic regimes, depending on how large the patch length $l$ is compared to the average superconducting coherence length $\xi$. If $l$ is less than $\xi$, the quasiparticle response of the superconductor to this order parameter profile is remarkably spatially uniform. In this regime, the LDOS from site to site does not vary appreciably. The order-parameter proxy observables show only very weak variations, and the distributions of both of these quantities are very narrow compared with that of the underlying $d$-wave order parameter. The spectra are effectively averaged over a coherence volume $\xi^{2}$, so the underlying atomic-scale variations in the order parameter are smoothed over in the quasiparticle response. If, on the other hand, $l$ is around the same magnitude as $\xi$, strikingly different behavior is observed which matches that seen in the superconducting cuprates. The LDOS at low energies is very uniform throughout the system, and deviations from uniformity only arise after a threshold energy is exceeded, beyond which the LDOS becomes strongly spatially fluctuating. We find that the kink energy demarcating the homogeneous-inhomogeneous crossover is generally set by the smallest set of gaps in the system, something that is also seen in experiment. In this regime, the statistical correlation between the local spectral gap and the height of the coherence peak is negative, similar to what is seen in STS experiments. Finally, in the regime where $l$ is larger than $\xi$, the behavior of the LDOS at a site where the local $d$-wave order parameter has a value $\Delta$ gradually reverts to that of a bulk $d$-wave superconductor with the same value of the order parameter. In this regime, the statistical correlations among the various observables also approach that expected from a bulk superconductor, with the proxy observables becoming very closely matched to the true order parameter. The locations of the kinks get pushed down all the way to zero energy with increasing $l$.
@@
We discuss in detail some ramifications of our results to the interpretation of STS experiments. In particular, we argue that in the cuprate superconductors, the behavior of the LDOS spectra and the spectral-gap distributions throughout much of the doping range are consistent with order-parameter patch sizes being in the $l\approx\xi$ regime. As a consequence, both the low-energy homogeneous $d$-wave state and the high-energy inhomogeneous state (but not the broken-symmetry density-wave and nematic states) can be explained within the same mean-field framework. We then reexamine the standard interpretation of the spectral gap as being a direct measure of the true superconducting order parameter and point out that in the $\l \approx\xi$ regime relevant to the cuprates, the width of the spectral-gap distribution vastly underestimates the width of the distribution of the underlying $d$-wave order parameter. It is then entirely plausible even that deep within the superconducting state of the cuprates, there can exist patches where the order parameter is very small or even zero, but proximity coupling of these with the surrounding large-order-parameter regions results in a low-energy homogeneous state within these regions that exhibits a finite spectral gap. We also consider estimates of the so-called “nodal gap” [12] and find that much of the observed “two-gap” phenomenology at low temperatures can be explained even within the inhomogeneous $d$-wave model to a surprisingly decent extent, with the dichotomy between the low-energy and high-energy states playing a crucial role. In addition, we discuss what our results mean for estimates of the local inelastic scattering rate in underdoped cuprates from STS data [10, 14]; we find that these estimates are likely an overestimate of the true inelastic scattering rate as these do not invoke the inhomogeneity of the pairing in the estimation process. Finally, we comment on the relation of order-parameter inhomogeneity to the hitherto unsolved phenomenon of QPI extinction [28, 29, 30] and suggest that inhomogeneity is of fundamental importance to this phenomenon in the first place, although we postpone a thorough discussion of this to a future paper.